OpenStudy (adrianna.gongora):

Simplify the square root of -121. A) –11i B) 11i C) –11 D) 11

OpenStudy (mathstudent55):

What is \(\sqrt{-1} \) ?

OpenStudy (adrianna.gongora):

I have no Idea I thought you couldn't square root a negative number.

OpenStudy (mathstudent55):

Using real numbers, you are correct. The square root of a negative number has no meaning.

OpenStudy (adrianna.gongora):

So then why did it give me options to choose from?

OpenStudy (adrianna.gongora):

The answer has to be one of these. A) –11i B) 11i C) –11 D) 11

OpenStudy (mathstudent55):

Let's go back to basics. The square root is the opposite operation of the square, the exponent 2. For example, \(5^2 = 25\). The opposite operation is the square root. \(\sqrt{25} = 5\)

OpenStudy (adrianna.gongora):

So then, \[\sqrt{-121}=11 ?\]

OpenStudy (mathstudent55):

Since a positive number times a positive number = positive And also a negative number times a negative number = positive, there is no numner you can multiply by itself and get a negative number.

OpenStudy (mathstudent55):

Since square root is the opposite of squaring, for real numbers, there does not exist a square root of a negative number.

OpenStudy (adrianna.gongora):

What does A and B mean then?

OpenStudy (mathstudent55):

That is why mathematicians invented imaginary numbers. The invention of imaginary numbers allows you to take the square root of a negative number.

OpenStudy (mathstudent55):

The basic idea with imaginary numbers is this definition: \(\sqrt{-1} = i\)

OpenStudy (adrianna.gongora):

So then it would be A) -11i

OpenStudy (mathstudent55):

Since squaring is the opposite operation of square root, then it follows that is \(\sqrt{-1} = i\), then \(i^2 = -1\)

OpenStudy (mathstudent55):

No, but we're almost there.

OpenStudy (adrianna.gongora):

Then B? 11i

OpenStudy (mathstudent55):

Keep in mind this basic idea of imaginary numbers: \(\sqrt{-1} = i\)

OpenStudy (adrianna.gongora):

Then it's B) 11i

OpenStudy (mathstudent55):

Now we deal with your problem. \(\sqrt{-121} = \sqrt{121 \times (-1)} = \sqrt{121} \times \sqrt{-1} = 11 \times i = 11i \)

OpenStudy (mathstudent55):

You are correct. It is choice B. The important thing is that you understand why it is 11i. I showed it to you step by step above.

OpenStudy (adrianna.gongora):

Thank you :) can you help me with a few more problems?

OpenStudy (mathstudent55):

Examples using \(\sqrt{-1} = i\) \(\sqrt{-4} = \sqrt{4} \times \sqrt{-1} = 2i\) \(-\sqrt{-144} = -\sqrt{144 \times (-1)} = -\sqrt{144}\times \sqrt{-1} = -12i\)

OpenStudy (mathstudent55):

You're welcome, and yes.

OpenStudy (adrianna.gongora):

It's number 2 I'm entirely lost.

OpenStudy (mathstudent55):

Please make a new post for each new question.

OpenStudy (adrianna.gongora):

Okay, one second.